OpenStudy (anonymous):

Find the radiance of convergence. sum_{n=1}^{infty} e^{-n^2}x^{n} is the radius 0? I used the root test and the one of the limit goes to 0..

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} e^{-n^2}x^{n}\]

OpenStudy (auctoratrox):

the radius of convergence of the infinite series is 1 by the root test.

OpenStudy (auctoratrox):

do you understand?

OpenStudy (anonymous):

abs(e^-n x)<1 then absx<inf R=inf

OpenStudy (anonymous):

for all x converge

OpenStudy (anonymous):

\[ \sum_{n=1}^{\infty} e^{-n^2}x^{n} \]By ratio test: \[ \frac{e^{-(n+1)^2x^{n+1}}}{x^ne^{-n^2}} = x \frac{e^{-(n+1)^2}}{e^{-n^2}} = 0\]So R = infinity, no?

OpenStudy (anonymous):

you shold test root thats beter i did

OpenStudy (anonymous):

your prove is said r=inf also

OpenStudy (anonymous):

so can I assume that whenever the limit goes to 0, R = infinity?

OpenStudy (anonymous):

I agree, but both yield the same result, after computing the limit, we will reach 0 < 1, which is true and shows that R = infinity. And, yeah, I was writing, didn't see your post ;-)

OpenStudy (anonymous):

@Dumboy You should think it like this: We got 0 and we want that the limit < 1 in order for the series to converge. For which values of x is 0 < 1? All of them, so R = infinity. So, yeah, if the limit goes to 0, you will get your answer as R = inf.

OpenStudy (anonymous):

it means for all real x you have 0<1 a true

OpenStudy (auctoratrox):

mahmit had it correct in his first post, and e^-n will go to 1 as x goes to infinity, so we get abs value of x is less than 1, so the radius is 1.

OpenStudy (anonymous):

there is no limit for x cause x is disappeared

OpenStudy (auctoratrox):

how did x disappear?

OpenStudy (anonymous):

dont worry just believe!

OpenStudy (auctoratrox):

using the root test leaves the abs. value of x is less than 1

OpenStudy (anonymous):

yeas are the same limit

OpenStudy (auctoratrox):

but the radius of convergence can't be infinity unless we do the limiting process and end up with a number that is less than 1, but we don't, we get x, correct?

OpenStudy (anonymous):

No, we get 0 :-)

OpenStudy (anonymous):

for all x correct means x belonging to all real

OpenStudy (auctoratrox):

and we do get 0 and 0,1 so r = infinity, ok

OpenStudy (anonymous):

ok E.x<1 then x<1/E and E go to 0 so x<inf

OpenStudy (anonymous):

R = inf